Grade 11Mechanics

A ball falls freely from a height of 1m on ground and rebounds to a height of 0.8m. Find coefficient of restitution.

Simran
10 Years agoGrade 11

2 Answers

arun
10 Years ago
let velocity of approach be \vartheta
then from third equation of motion
\vartheta^2 = u^2 +2as
and in his case u=0 m/s
a=g
s= 1 metres
\Rightarrow \vartheta =\sqrt{2g}
now let velocity of seperation is v
again using third equation of motion
v^2 = u^2 +2as
and in this case 
 u=0 m/s
a=g
s = 0.8 metres
v=\sqrt{1.6g}
now e =v\div \vartheta
e= \sqrt{1.6\div 2}
e= \sqrt{0.8}
J bat
8 Years ago
The coefficient of restitution is given bye^n = ( hn/h)^1/2 .: for 1 collision n=1 ,hn = 0.8, h=1e=(0.8)^1/2